OpenStudy (anonymous):

Which quadratic equation has solutions x = 1/2 and x = 2/3?

OpenStudy (anonymous):

x=1/2 2x=1 2x-1=0 (2x-1) x=2/3 3x=2 3x-2=0 (3x-2) (2x-1)(3x-2) -->foil 6x² -3x -4x +2 6x²-7x+2=0

OpenStudy (anonymous):

Thanks!